airamchavez:

This object is composed of a cylinder of diameter 4 cm and height 14 cm on top of another cylinder of diameter 12 cm and height 4 cm. Determine the surface area of the object, to the nearest square centimeter.

Ashely:

do you possibly have a pic

surjithayer:

 surface areaof smaller cylinder\[=2\pi r1h+\pi (r1)2 ^2\] \[=2\pi (2)(14)+\pi(2)^2\] \[=56 \pi+4\pi \] \[=60 \pi\] \[surface~area~of~bigger~cylinder =2 \pi r2h+2 (\pi (r2)^2)-\pi (r1)^2\] \[=2 \pi (6)(4)+2 \pi (6)^2-\pi (2)^2\] \[=48 \pi+72 \pi-4 \pi\] \[=116 \pi\] total surface area=?